ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

8. Analysis

1991 MS Classification: 33E10, 43A62, 43A10 Connett, William, University of Missouri-St. Louis, USA Measure algebras that have oblate spheroidal wave functions as characters The recently discovered product formula for the oblate spheroidal wave functions can be used to describe a convolution structure for a number of measure algebras, some of which are historically important examples of non-compact hypergroups, some of which are new hypergroups, and some of which are more general measure algebras (for example, signed hypergroups).